28. Il Peregrine falcons frequently grab prey birds from the air, as BI0 in Example 9.10. Sometimes they strike at high enough speeds that the force of the impact disables prey birds. A 480 g per- egrine falcon high in the sky spies a 240 g pigeon some dis- tance below. The falcon slows to a near stop, then goes into ve-called a stoop-and picks up speed as she falls. The n reaches a vertical speed of 45 m/s before striking the pigeon, which we can assume is stationary. The falcon strikes the pigeon and grabs it in her talons. The collision between the falco birds lasts 0.015 s. a. What is the final speed of the falcon and pigeon? b. What is the average force on the pigeon during the impact?
Answer #1

Given,

m1 = 480 g = 0.48 kg ; m2 = 240 g = 0.24 kg ; v1 = 45 m/s ; v2 = 0 ; t = 0.015 s

a)The birds will move with a common speed after falcon grabs the pigeon. Let this ve vf.

From conservation of momentum

m1v1 + m2v2 = (m1 + m2)vf

sinc2 v2 = 0

vf = m1v1/(m1 + m2)

vf = 0.48 x 45/(0.48 + 0.24) = 30 m/s

Hence, vf = 30 m/s

b)We know that, impulse is:

I = m (v2 - v1)

also, I = F t

m (v2 - v1) = F t

F = m (v2 - v1)/t

F = 0.48 (30 - 45)/0.015 = -480 N

Hence, F = -480 N

(you can ignore negative sign for magnitude only)